# Reparameterized LLM Training via Orthogonal Equivalence Transformation

### v3

@@ 5.2 Ablation Studies and Empirical Analyses
Initialization schemes. We empirically compare different random initialization schemes for POET, including two commonly used ones (standard Gaussian, Xavier [17]) and two proposed ones (uniform spectrum, normalized Gaussian). Specifically, we use fully stochastic POET with $b$=1/2 to train Llama-60M on 30B tokens and report the validation perplexity in Table 3. Results show that the normalized initialization will lead to the best final performance, and we stick to it as a default choice. Interestingly, uniform spectrum initialization performs poorly. This suggests a trade-off between preserving good weight spectral properties and achieving strong expressiveness. it may limit its expressiveness. Finding the optimal singular value structure for weights remains an important open problem.
@@
Merge-then-reinitialize frequency. The proposed merge-then-reinitialize trick allows POET to train only a small fraction of the large orthogonal matrices $\bm{R}$ and $\bm{P}$ per iteration, significantly reducing GPU memory usage. However, this trick also introduces a reinitialization frequency hyperparameter $T_{m}$, which determines how often the orthogonal matrix is merged and reset to the identity. The index set in POET-FS and the permutation matrix in POET-BS are also resampled at each reinitialization. Therefore, it is quite important to understand how this hyperparameter $T_{m}$ affects performance. Following the previous initialization experiment, we use POET-FS with $b$=1/2 to train Llama-60M on 30B tokens. We vary the reinitialization frequency from $5$ to $1600$ and report the validation perplexity in Table 4. Results show that both $200$ and $400$ perform well. Therefore, we set $T_{m}=400$ in all experiments by default.
@@
Neumann series approximation. CNP approximates the matrix inverse using a Neumann series. As the number of Neumann terms directly influences the approximation quality, understanding its impact on model performance is essential. To this end, we evaluate how varying the number of Neumann terms affects performance, using POET-FS with $b$ = 1/2 to train LLaMA-130M. Results in Table 5 show that increasing the number of Neumann terms generally improves validation perplexity. However, this also leads to slower training. Moreover, Using only $1$ Neumann term ($k=1$) leads to training divergence, highlighting the critical role of maintaining orthogonality. To balance overhead and performance, we find that using $5$ Neumann terms is a good trade-off.
@@
Additionally, it is important to evaluate the accuracy of the Neumann approximation to understand how the number of Neumann terms affects the preservation of orthogonality. The orthogonal approximation error is defined by $e_{\text{orth}}={\|RR^{T}-I\|_{F}}/{\|I\|_{F}}$. We randomly select a weight matrix and compute the approximation error of two orthogonal matrices $\bm{R}$ and $\bm{P}$ that correspond to it. For clarity, we only visualize the error in the initial 1000 training steps in Figure 10. We can observe that, with more Neumann terms, the orthogonal approximation error is indeed lower. We also note that the merge-then-reinitialize trick periodically resets the error.
Method # Params XSum CNN/DailyMail
LoRA ($r$=32) 17.30M 43.38 / 20.20 / 35.25 43.17 / 20.31 / 29.72
OFT ($b$=64) 8.52M 44.12 / 20.96 / 36.01 44.08 / 21.02 / 30.68
Full FT 406.29M 45.14 / 22.27 / 37.25 44.16 / 21.28 / 40.90
POET(FS,$b$=1/2) 144.57M 45.23 / 22.41 / 37.28 44.27 / 21.29 / 41.02
POET for finetuning. To demonstrate the applicability of POET to general finetuning tasks, we apply it to finetune a BART-large model [41] on the NLP task of text summarization. Specifically, we evaluate POET on the XSum [61] and CNN/DailyMail [24] datasets, reporting ROUGE-1/2/L scores in Table 6. We note that both LoRA and OFT are designed solely for parameter-efficient finetuning and are not applicable to pretraining. Our goal here is to demonstrate that POET is also effective as a finetuning method. For consistency, we use the same configuration as in the pretraining setup, resulting in a higher parameter count. Experimental results show that POET not only supports finetuning effectively but also outperforms both full-model finetuning and parameter-efficient methods.
6 Related Work and Concluding Remarks
@@ A.3 Empirical Observations
Phase I: conical-shell searching rotates each orthogonal matrix $R$ and $P$ smoothly away from the identity while preserving their singular values, so the cosine similarity between transformed and initial weight vectors falls from $1$ to $\approx 0.6$; this provides a spectrally well-conditioned “cone” in which learning can proceed safely. this phase serves the role of “spectral warm-up”. By plotting the cosine similarity of any one layer, we always see the same graceful slide towards 0.60.60.60.60.650.650.650.65, independent of model size, layer type, or whether you train with fully-stochastic or block-stochastic SPO. This phase carves out the thin “shell” in which subsequent learning lives.
@@ Appendix C Guarantees of Weight Spectrum under POET
For standard Gaussian initialization where each element of the weight matrix $\bm{W}\in d\times n$ is sampled with a normal distribution, we have the following standard results [70, 8]: {mdframed}
[…]
which gives spectrum guarantees for weight matrices generated by the standard Gaussian initialization.
@@
Let $\tilde{\bm{v}}_{1},\cdots,\tilde{\bm{v}}_{n}\in\mathbb{R}^{d}$ be i.i.d. random vectors where each element follows the Gaussian distribution with mean 00 and variance $1$. Then $\bm{v}_{1}=\frac{\tilde{\bm{v}}_{1}}{\|\tilde{\bm{v}}_{1}\|_{2}},\cdots,\bm{v}_{n}=\frac{\tilde{\bm{v}}_{n}}{\|\tilde{\bm{v}}_{n}\|_{2}}$ are uniformly distributed on the unit hypersphere $\mathbb{S}^{d-1}$. If the ratio $\frac{n}{d}$ converges to a constant $\lambda\in(0,1)$, asymptotically we have for $\bm{W}=\{\bm{v}_{1},\cdots,\bm{v}_{n}\}\in\mathbb{R}^{d\times n}$:
@@
$\lim_{n\rightarrow\infty}\sigma_{\max}(\tilde{\bm{W}})$ $=\sqrt{d}+\sqrt{\lambda d}$ (21)
$\lim_{n\rightarrow\infty}\sigma_{\min}(\tilde{\bm{W}})$ $=\sqrt{d}-\sqrt{\lambda d}$
@@
[…]
@@
which can be written to the following results: {mdframed}
[…]
which shows that under our proposed normalized Gaussian initialization, the maximal and minimal singular values are well bounded by a constant that is only dependent on the size of weight matrix. These results justify the effectiveness of our proposed normalized Gaussian initialization in POET.
@@ Appendix E Experimental Details
We employed the AdamW optimizer [55] for all our training runs. The specific hyperparameters used for each experiment are detailed in the Table 8 and Table 9 referenced below. We use the consine learning rate scheduler with the minimum learning ratio of $0.01$. We use the number of warmup steps of 00, weight decay of $0.01$ and gradient clipping of $0.1$. For the AdamW baseline, we report results for the optimal learning rate from $[1{\times}10^{-2},5{\times}10^{-3},1{\times}10^{-3},5{\times}10^{-4},1{\times}10^{-4},5{\times}10^{-5},1{\times}10^{-5}]$. After each merge-then-reinitalize step, we additionally increase the gradient clipping for 10 training steps to improve training stability.
@@
Our work utilized the Hugging Face Transformers [^2] https://github.com/huggingface/transformers code base to construct the Llama model for pertaining, which is under the Apache 2.0 license. The specific layer setups for the different scaled Llama models are summarized in Table 7. Note, the intermediate dimension of the Feed-Forward Network (FFN) has been slightly modified for the POET-BS, compared to the configs in [81], because the linear layer dimensions have to be divisible by the POET-BS block size $b$.
@@
We use the Colossal Clean Crawled Corpus (C4) dataset [66] for pertaining. The C4 data is a large-scale, meticulously cleaned version of Common Crawl’s web crawl corpus. It was originally introduced for training the Text-to-Text Transfer Transformer (T5) model and has since become a standard pre-training dataset for testing training algorithms for pre-training large language models. The dataset is released under the ODC-BY license.
@@ Appendix F Implementation and CUDA Acceleration
To enable efficient POET training, we implement the Cayley–Neumann parameterization. To reduce memory usage, we leverage the structure of the skew-symmetric matrix $\bm{Q}\in\mathbb{R}^{n\times n}$, where the diagonal entries are zero ($Q_{ii}=0$) and off-diagonal elements satisfy $Q_{ij}=-Q_{ji}$. This structure allows us to store only the upper triangular part of $\bm{Q}$ as a vector, reducing the number of trainable parameters from $n^{2}$ to $n(n-1)/2$. During the forward pass, $\bm{Q}$ is reconstructed on-the-fly using a specialized CUDA kernel, significantly accelerating this process. In addition, the Neumann approximation removes the need for costly and numerically unstable matrix inversion, offering further computational gains. Overall, training a 1.3B LLaMA model on a single H100 8-GPU node yields a 3.8× speedup over the baseline (i.e., native implementation). Table 10 summarizes the contribution of each component to the overall training time.
@@
Neumann approximation 1.5$\times$
Skew-symmetric CUDA kernel 1.3$\times$
Total 3.8$\times$
@@ Appendix H Weight Update Evenness of Different POET Variants
To understand the higher parameter efficiency of POET-BS compared to POET-FS, we employ a toy example to visualize their different weight update mechanisms by counting the total number of updates for each element of the weight matrix. The visualization results are given in Figure 15 and Figure 16. Specifically, in this experiment, a 64$\times$64 matrix was randomly initialized and trained for 100 steps under various POET-BS and POET-FS configurations. The merge-then-reinitialize trick is performed at each iteration, and the same set of weight elements was effectively updated between two successive merge-then-reinitialize operations. For each weight element, we compute its total number of update in these 100 steps.
@@ Appendix I Training Dynamics of Singular Values
Figure 17: Training dynamics of the singular values of weight matrices within Blocks 0–1 (the $i$-th row represents Block $i$) of a 60M Llama Transformer trained with AdamW.
Figure 18: Training dynamics of the singular values of weight matrices within Blocks 2–4 (the $i$-th row represents Block $i$) of a 60M Llama Transformer trained with AdamW.
Figure 19: Training dynamics of the singular values of weight matrices within Blocks 5–7 (the $i$-th row represents Block $i$) of a 60M Llama Transformer trained with AdamW.
@@
In this ablation study, we evaluate the approximation error of the orthogonal matrices $\bm{R}\in\mathbb{R}^{m\times m}$ and $\bm{P}\in\mathbb{R}^{n\times n}$ across all linear layers in Block 0 of a 130M LLaMA model trained with POET-FS ($b=1/2$) for 10,000 steps. Figure 23 and Figure 24 show the approximation error over the first 1,000 steps. Since the error difference between $k=4$ and $k=5$ was negligible, we used $k=4$ for better computational efficiency. Empirically, while $k=2$ or $k=3$ suffices for smaller LLaMA models, larger $k$ values are needed to avoid training divergence caused by exploding gradients due to approximation error.
@@
Additionally, Figure 25 shows the orthogonality approximation error of Neumann series with different $k$ over the first 10,000 training steps, illustrating how it decreases as training progresses. We observe a general downward trend in approximation error, indicating improved approximation over time. The results also suggest that using too few Neumann series terms (e.g., $k=1$) can lead to training divergence in POET.

### v4

@@ 5.3 Ablation Studies and Empirical Analyses
Initialization schemes. We empirically compare different random initialization schemes for POET, including two commonly used ones (standard Gaussian, Xavier [17]) and two proposed ones (uniform spectrum, normalized Gaussian). Specifically, we use fully stochastic POET with bb=1/2 to train Llama-60M on 30B tokens and report the validation perplexity in Table 5. Results show that the normalized initialization will lead to the best final performance, and we stick to it as a default choice. Interestingly, uniform spectrum initialization performs poorly. This suggests a trade-off between preserving good weight spectral properties and achieving strong expressiveness. it may limit its expressiveness. Finding the optimal singular value structure for weights remains an important open problem.
@@
Merge-then-reinitialize frequency. The proposed merge-then-reinitialize trick allows POET to train only a small fraction of the large orthogonal matrices $\bm{R}$ and $\bm{P}$ per iteration, significantly reducing GPU memory usage. However, this trick also introduces a reinitialization frequency hyperparameter $T_{m}$, which determines how often the orthogonal matrix is merged and reset to the identity. The index set in POET-FS and the permutation matrix in POET-BS are also resampled at each reinitialization. Therefore, it is quite important to understand how this hyperparameter $T_{m}$ affects performance. Following the previous initialization experiment, we use POET-FS with bb=1/2 to train Llama-60M on 30B tokens. We vary the reinitialization frequency from $5$ to $1600$ and report the validation perplexity in Table 6. Results show that both $200$ and $400$ perform well. Therefore, we set $T_{m}=400$ in all experiments by default.
@@
Neumann series approximation. CNP approximates the matrix inverse using a Neumann series. As the number of Neumann terms directly influences the approximation quality, understanding its impact on model performance is essential. To this end, we evaluate how varying the number of Neumann terms affects performance, using POET-FS with $b$ = 1/2 to train LLaMA-130M. Results in Table 7 show that increasing the number of Neumann terms generally improves validation perplexity. However, this also leads to slower training. Moreover, dropping all Neumann terms ($k=0$) leads to training divergence, highlighting the critical role of maintaining orthogonality. To balance overhead and performance, we find that $3$ Neumann terms ($k=3$) reach a good trade-off.
@@
Additionally, we evaluate the accuracy of the Neumann approximation to understand how the number of Neumann terms affects the orthogonality. The orthogonal approximation error is defined by $e_{\text{orth}}={\|\bm{R}\bm{R}^{T}-\bm{I}\|_{F}}/{\|\bm{I}\|_{F}}$. We randomly select a weight matrix and compute the approximation error of corresponding orthogonal matrices $\bm{R}$ and $\bm{P}$. For clarity, we visualize the error in the initial 1000 training steps in Figure 10. We observe that, with more Neumann terms, the orthogonal approximation error is indeed lower. Moreover, the merge-then-reinitialize trick can periodically reset the error. More results are given in Appendix L
6 Related Work, Concluding Remarks and Acknowledgement
@@
Acknowledgement. The authors would like to sincerely thank Lixin Liu, Han Shi, Gege Gao, Zhen Liu and many colleagues at Max Planck Institute for Intelligent Systems for many helpful suggestions. Additionally, the authors also sincerely thank all the anonymous NeurIPS reviewers for their constructive suggestions that greatly improved the quality of our work.
@@ A.3 Empirical Observations
Phase I: conical-shell searching rotates each orthogonal matrix $R$ and $P$ smoothly away from the identity while preserving their singular values, so the cosine similarity between transformed and initial weight vectors falls from $1$ to $\approx 0.6$; this provides a spectrally well-conditioned “cone” in which learning can proceed safely. this phase serves the role of “spectral warm-up”. By plotting the cosine similarity of any one layer, we always see the same graceful slide towards 0.60.60.650.65, independent of model size, layer type, or whether you train with fully-stochastic or block-stochastic SPO. This phase carves out the thin “shell” in which subsequent learning lives.
@@ Appendix C Guarantees of Weight Spectrum under POET
For standard Gaussian initialization where each element of the weight matrix $\bm{W}\in d\times n$ is sampled with a normal distribution, we have the following standard results [70, 8]: $\frac{1}{\sqrt{d}}\sigma_{\max}({\bm{W}})$ $\xrightarrow[n\rightarrow\infty]{\text{a.s.}}1+\sqrt{\lambda}$ (18) $\frac{1}{\sqrt{d}}\sigma_{\min}({\bm{W}})$ $\xrightarrow[n\rightarrow\infty]{\text{a.s.}}1-\sqrt{\lambda}$ which gives spectrum guarantees for weight matrices generated by the standard Gaussian initialization.
@@
Let $\tilde{\bm{v}}_{1},\cdots,\tilde{\bm{v}}_{n}\in\mathbb{R}^{d}$ be i.i.d. random vectors where each element follows the Gaussian distribution with mean 0 and variance $1$. Then $\bm{v}_{1}=\frac{\tilde{\bm{v}}_{1}}{\|\tilde{\bm{v}}_{1}\|_{2}},\cdots,\bm{v}_{n}=\frac{\tilde{\bm{v}}_{n}}{\|\tilde{\bm{v}}_{n}\|_{2}}$ are uniformly distributed on the unit hypersphere $\mathbb{S}^{d-1}$. If the ratio $\frac{n}{d}$ converges to a constant $\lambda\in(0,1)$, asymptotically we have for $\bm{W}=\{\bm{v}_{1},\cdots,\bm{v}_{n}\}\in\mathbb{R}^{d\times n}$:
@@
[…]
@@
$\begin{aligned} \lim_{n\rightarrow\infty}\sigma_{\max}(\bm{W})&=\lim_{n\rightarrow\infty}\sigma_{\max}(\tilde{\bm{W}}\cdot\bm{Q})\\ \lim_{n\rightarrow\infty}\sigma_{\min}(\bm{W})&=\lim_{n\rightarrow\infty}\sigma_{\min}(\tilde{\bm{W}}\cdot\bm{Q})\end{aligned}.$ (23)
@@
which can be written to the following results: $\sigma_{\max}({\bm{W}})$ $\xrightarrow[n\rightarrow\infty]{\text{a.s.}}1+\sqrt{\lambda}$ (36) $\sigma_{\min}({\bm{W}})$ $\xrightarrow[n\rightarrow\infty]{\text{a.s.}}1-\sqrt{\lambda}$ which shows that under our proposed normalized Gaussian initialization, the maximal and minimal singular values are well bounded by a constant that is only dependent on the size of weight matrix. These results justify the effectiveness of our proposed normalized Gaussian initialization in POET.
@@ Appendix E Experimental Details
We employed the AdamW optimizer [55] for all our training runs. The specific hyperparameters used for each experiment are detailed in the Table 9 and Table 10 referenced below. We use the consine learning rate scheduler with the minimum learning ratio of $0.01$. We use the number of warmup steps of 0, weight decay of $0.01$ and gradient clipping of $0.1$. For the AdamW baseline, we report results for the optimal learning rate from $1{\times}10^{-2},5{\times}10^{-3},1{\times}10^{-3},5{\times}10^{-4},1{\times}10^{-4},5{\times}10^{-5},1{\times}10^{-5}]$. After each merge-then-reinitalize step, we additionally increase the gradient clipping for 10 training steps to improve training stability.
@@
Our work utilized the Hugging Face Transformers [^2] https://github.com/huggingface/transformers code base to construct the Llama model for pretraining, which is under the Apache 2.0 license. The specific layer setups for the different scaled Llama models are summarized in Table 8. Note, the intermediate dimension of the Feed-Forward Network (FFN) has been slightly modified for the POET-BS, compared to the configs in [82], because the linear layer dimensions have to be divisible by the POET-BS block size $b$.
@@
We use the Colossal Clean Crawled Corpus (C4) dataset [66] for pretraining. The C4 data is a large-scale, meticulously cleaned version of Common Crawl’s web crawl corpus. It was originally introduced for training the Text-to-Text Transfer Transformer (T5) model and has since become a standard pre-training dataset for testing training algorithms for pre-training large language models. The dataset is released under the ODC-BY license.
@@ Appendix F Implementation and CUDA Acceleration
To enable efficient POET training, we implement the Cayley–Neumann parameterization. To reduce memory usage, we leverage the structure of the skew-symmetric matrix $\bm{Q}\in\mathbb{R}^{n\times n}$, where the diagonal entries are zero ($Q_{ii}=0$) and off-diagonal elements satisfy $Q_{ij}=-Q_{ji}$. This structure allows us to store only the upper triangular part of $\bm{Q}$ as a vector, reducing the number of trainable parameters from $n^{2}$ to n(n1)/2n(n-1)/2. During the forward pass, $\bm{Q}$ is reconstructed on-the-fly using a specialized CUDA kernel, significantly accelerating this process. In addition, the Neumann approximation removes the need for costly and numerically unstable matrix inversion, offering further computational gains. Overall, training a 1.3B LLaMA model on a single H100 8-GPU node yields a 3.8× speedup over the baseline (i.e., native implementation). Table 11 summarizes the contribution of each component to the overall training time.
@@
[…]
@@ Appendix H Weight Update Evenness of Different POET Variants
To understand the higher parameter efficiency of POET-BS compared to POET-FS, we employ a toy example to visualize their different weight update mechanisms by counting the total number of updates for each element of the weight matrix. The visualization results are given in Figure 15 and Figure 16. Specifically, in this experiment, a $\times64$ matrix was randomly initialized and trained for 100 steps under various POET-BS and POET-FS configurations. The merge-then-reinitialize trick is performed at each iteration, and the same set of weight elements was effectively updated between two successive merge-then-reinitialize operations. For each weight element, we compute its total number of update in these 100 steps.
@@ Appendix I Training Dynamics of Singular Values
Figure 17: Training dynamics of the singular values of weight matrices within Blocks 0–1 (the $i$-th row represents Block ii) of a 60M Llama Transformer trained with AdamW.
Figure 18: Training dynamics of the singular values of weight matrices within Blocks 2–4 (the $i$-th row represents Block ii) of a 60M Llama Transformer trained with AdamW.
Figure 19: Training dynamics of the singular values of weight matrices within Blocks 5–7 (the $i$-th row represents Block ii) of a 60M Llama Transformer trained with AdamW.
@@
In this ablation study, we evaluate the approximation error of the orthogonal matrices $\bm{R}\in\mathbb{R}^{m\times m}$ and $\bm{P}\in\mathbb{R}^{n\times n}$ across all linear layers in Block 0 of a 130M LLaMA model trained with POET-FS ($b=1/2$) for 10,000 steps. Figure 23 and Figure 24 show the approximation error over the first 1,000 steps. Since the error difference between $k=3$ and $k=4$ was negligible, we used $k=3$ for better computational efficiency. Empirically, while $k=1$ or $k=2$ suffices for smaller LLaMA models, larger $k$ values are needed to avoid training divergence caused by exploding gradients due to approximation error.
@@
Additionally, Figure 25 shows the orthogonality approximation error of Neumann series with different $k$ over the first 10,000 training steps, illustrating how it decreases as training progresses. We observe a general downward trend in approximation error, indicating improved approximation over time. The results also suggest that using too few Neumann series terms can lead to training divergence in POET.
@@
Appendix L More Results of POET as a Finetuning Method
To demonstrate the applicability of POET to general finetuning tasks, we apply it to finetune a BART-large model [41] on the NLP task of text summarization. Specifically, we evaluate POET on the XSum [61] and CNN/DailyMail [24] datasets, reporting ROUGE-1/2/L scores in Table 12. We note that both LoRA and OFT are designed solely for parameter-efficient finetuning and are not applicable to pretraining. Our goal here is to demonstrate that POET is also effective as a finetuning method. For consistency, we use the same configuration as in the pretraining setup, resulting in a higher parameter count. Experimental results show that POET not only supports finetuning effectively but also outperforms both full-model finetuning and parameter-efficient methods.
Method # Params XSum CNN/DailyMail
LoRA (rr=32) 17.30M 43.38  /  20.20  /  35.25 43.17  /  20.31  /  29.72
OFT (bb=64) 8.52M 44.12  /  20.96  /  36.01 44.08  /  21.02  /  30.68
Full FT 406.29M 45.14  /  22.27  /  37.25 44.16  /  21.28  /  40.90
POET (FS,bb=1/2) 144.57M 45.23  /  22.41  /  37.28 44.27  /  21.29  /  41.02